a dice is rolled the probability of getting
  1. a number greater than 2
  2. a odd number                        are they independent events if yes i cant see how practically
 

Grade:12

3 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

Sample space, S is
S = {1, 2, 3, 4, 5, 6}
(i) Let E be the event of occuring a number on die greater than 2
E = {3, 4, 5, 6}
Probability = 4/6 = 2/3

(ii) Let E be the event of occuring odd number on die
E = {1, 3, 5}
Probability = 3/6 = 1/2

rohan
9 years ago
no, i asked whether they are independent and how can this happen that the two events in single dice is independent
Jitender Singh
9 years ago
Two events are independent if the occurence of one event does not effect the probability of other.
Lets say 1st event to be A and 2nd to be B. Let C to be intersection of A and B. A and B are independent if
P(C) = P(A)P(B)
C = {3, 5}
P(C) = 2/6 = 1/3
P(A)P(B) = 2/3 . ½ = 1/3
Hence, A & B are independent.